A consumer protection agent wants to test a paint manufactures claim that the average drying time of his new fast-drying paiPLEASE SHOW ALL WORK THANKS
Answer #1

X ~ N ( µ = 20 , σ = 2.4 )
P ( X > 20.5 ) = 1 - P ( X < 20.5 )
Standardizing the value
Z = ( X - µ ) / ( σ / √(n))
Z = ( 20.5 - 20 ) / ( 2.4 / √ ( 36 ) )
Z = 1.25
P ( ( X - µ ) / ( σ / √ (n)) > ( 20.5 - 20 ) / ( 2.4 / √(36) )
P ( Z > 1.25 )
P ( X̅ > 20.5 ) = 1 - P ( Z < 1.25 )
P ( X̅ > 20.5 ) = 1 - 0.8944
P ( X̅ > 20.5 ) = 0.1056

P ( Type I error ) α = 0.1056

X ~ N ( µ = 21 , σ = 2.4 )
P ( X < 20.5 )
Standardizing the value
Z = ( X - µ ) / (σ/√(n)
Z = ( 20.5 - 21 ) / ( 2.4 / √36 )
Z = -1.25
P ( ( X - µ ) / ( σ/√(n)) < ( 20.5 - 21 ) / ( 2.4 / √(36) )
P ( X < 20.5 ) = P ( Z < -1.25 )
P ( X̅ < 20.5 ) = 0.1056

P ( Type II error ) ß = 0.1056

Option a is correct.